Electronic Notes in Theoretical Computer Science XX (2000)

URL: http://www.elsevier.nl/locate/entcs/volumeXX.html

An Example Paper

My Name^{1,2}

My Department My University My City, My Country

My Co-author^{3}

My Co-author’s Department My Co-author’s University My Co-author’s City, My Co-author’s Country

Abstract This is a short example to show the basics of using the ENTCS style macro files. Ample examples of how files should look may be found on the ENTCS Home Page

http://www.elsevier.nl/locate/entcs.

1 Introduction

This short note provides a guide to using the ENTCS macro package for preparing papers for publication in your conference Proceedings. The Proceedings may be printed and hard copies distributed to participants at the meeting; this is an option to Conference Organizers may choose to exercise. The Proceedings also will be a volume in the series Electronic Notes in Theoretical Computer Science (ENTCS), which is published under the auspices of Elsevier Science B. V., the publishers of Theoretical Computer Science.

The ENTCS macro package consists of two files:

entcs.cls, the basic style file, and entcsmacro.sty, a macro file containing the definitions of some of the theorem-

like environments and a few other tidbits.

1 Thanks to everyone who should be thanked

2 Email: myuserid@mydept.myinst.myedu
3 Email: couserid@codept.coinst.coedu

\begin{matrix}c \\ ◯ \end{matrix}2000 Published by Elsevier Science B. V.

My Lastname

Additional macro files can be added using \usepackage{...}. The file entcsmacro.sty must be included in the list, as is done at the start of the source file for this paper.

The formatting these style files impose should not be altered – the reason for using them is to attain a uniform format for all papers in the Proceedings of which your paper is a part.

The file instraut.dvi contains information about the macro package

elsart.sty and elsart12.sty, from which the files entcs.sty and entcs12.sty

are derived. Reading this file should answer most of the basic questions that might arise. It also gives some basic rules about using the macro package.

2 Frontmatter

The biggest difference between a “usual” L^{A}TEX style such as article.sty and the ENTCS package is that the ENTCS macro package requires the title, author’s name or names, abstract and “thanks” all to be included within the frontmatter environment. At the beginning of the source file for this paper, you’ll notice this. Also, you’ll notice that the usual \maketitle is absent; it no longer is needed. The ENTCS style package automatically generates the title, author’s name and address, and related material at the beginning of the paper.

The ENTCS macro package provides two alternatives to listing authors names and addresses. These are described in detail in the file instraut.dvi. Basically, listing each author and his or her address in turn, is the simplest method. But, if there are several authors and two or more share the same address (but not all authors are at this address), then the method of listing authors first, and then the addresses, and of referencing addresses to authors should be used.

Also, notice that acknowledgment of support (the contents of \thanks) should be done by a separate listing of \thanks[NSF]{To the NSF} with the optional argument – [NSF] – being used for \thanksref which is attached to those authors acknowledging such support. It is important that the \thanks not be included within the scope of \author{} or of \title{}, but it must be within the scope of the environment frontmatter.

More details about added terms such as \collab can be found in inst.dvi, if they are needed.

Also, notice that the command \lastname{My Lastname} has been included before the frontmatter begins. This command should contain the last names of the authors of the paper. If there are no more than three authors, then they should be listed with the word “and” between the last two; if more than three authors collaborated on the paper, then the first author only should be listed, together with \emph{et al}. This command creates the headline for each page after page 1.

Finally, please be sure to include an abstract for your paper.

3 Sectioning and Environments

Since ENTCS is published through the auspices of Elsevier Science B. V., their style files have been used to create the ENTCS macro package. Here’s a proof that this package is not much different than most of the ones one encounters:

Definition 3.1 A file is derived from another if it is obtained with only a few modifications from the original file.

Theorem 3.2 The file entcs.cls is derived from elsart.sty.

Proof. This is clear from the similarity of the output to the output from Elsevier’s style files. □

If one wants to start a proof with a descriptive word, such as “sketch”, then one can use the \begin{proof*}...\end{proof*} environment, as in

Proof (Sketch) This can be derived from simple observations. □

The main differences between the file entcs.cls and the ones created by Elsevier are the more precise format we use – Elsevier’s generic files are meant for preliminary editing, and more precise formatting is imposed using a macro file designed for the specific Elsevier journal in which the paper will eventually appear. The entcs.cls and entcsmacro.sty files format papers uniformly so that they all are easily recognizable as being from the series Electronic Notes in Theoretical Computer Science.

All of the usual features of L^{A}TEX are available with these style files – it is only the formatting that has been rigorously defined. Thus, one has available the sectioning commands \section,\subsection, \paragraph and \subparagraph. The numbering scheme used is one under which Theorem 1.2.3 is the third numbered item in second subsection of the first section of the paper. In order to facilitate cross-references, all of the named environments given below are numbered, and all use the same number scheme.

The file entcsmacro.sty contains additional information that is needed to typeset a paper. It also has the definitions of the AMS euler and blackboard bold fonts builtin. If you want to use symbols for the natural numbers, the reals, etc., then we prefer that you use the blackboard bold fonts, and not plain bold fonts. This is accomplished by using the \mathbb font, as in ℕ or ℝ.

The names of theorem-like environments are provided in entcsmacro.sty. With the exception of the environment Algorithm, the names of all of these are full name, rather than a shortened version. The environments provided and their names are

• \begin{theorem} ... \end{theorem} for Theorems,
• \begin{lemma} ... \end{lemma} for Lemmas,
• \begin{corollary} ... \end{corollary} for Corollaries,

• \begin{proposition} ... \end{proposition} for Propositions,
• \begin{criterion} ... \end{criterion} for Criteria,
• \begin{alg} ... \end{alg} for Algorithms,
• \begin{definition} ... \end{definition} for Definitions,
• \begin{conjecture} ... \end{conjecture} for Conjectures,
• \begin{example} ... \end{example} for Examples,
• \begin{problem} ... \end{problem} for Problems,
• \begin{remark} ... \end{remark} for Remarks,
• \begin{note} ... \end{note} for Notes,
• \begin{claim} ... \end{claim} for Claims,
• \begin{summary} ... \end{summary} for Summary,
• \begin{case} ... \end{case} for Cases, and
• \begin{ack} ... \end{ack} for Acknowledgements.

For example,

Algorithm 1 Step 1: Write the paper Step 2: Format it with the ENTCS macro package Step 3: Ship the whole thing to the Guest Editors

4 References and Cross-references

All the cross-referencing facilities of L^{A}TEX are supported, so one can use \ref{} and \cite{} for cross-references within the paper and for references to bibliographic items. As is done in this note, the References section 6 can

be composed with \begin{thebibliography}...\end{thebibliography}.

Alternatively, BibTEX can be used to compile the bibliography. Whichever one is used, the references are to be numbered consecutively, rather than by author-defined acronyms. Of course you can use your own acronyms for easy reference to each of the items in the bibliography, as has been done with the listing for this short note.

However, note that the references should not be started with a new \section command.

The package hyperref is automatically loaded by entcs.cls, and this makes all the cross-references within the document “active” when the file is viewed with xdvi. The format for including a link is simple: simply insert \href{URL} {text} where URL is the URL to which you want the link to point, and text is the text you want to be highlighted, which when clicked upon will bring up the desired web page. L^{A}TEX appears to have problems dealing with the crossreferences embedded within the frontmatter section, but the complaints appear to be harmless, and simply pushing L^{A}TEX to run through them appears to work. We are working on the problem – if some L^{A}TEXpert figures out how to cure it (other than the unacceptable method of disabling hyperref), we would appreciate hearing about it.

4.1 Particulars about .pdf files

In addition to PostScript^{\begin{matrix}c \\ ◯ \end{matrix}} files of final versions of papers appearing ENTCS, we also require that .pdf files be provided for publication online. A .pdf file is viewable by Adobe’s Acrobat^{\begin{matrix}c \\ ◯ \end{matrix}} viewer, which can be configured to load automatically within a browser. Viewing a properly formatted .pdf file with Acrobat^{\begin{matrix}c \\ ◯ \end{matrix}} allows the cross-references and links to URLs to be active. In fact, we have decided to utilize .pdf files in order to take better advantage of the web’s capabilities. There are two methods for producing .pdf files.

The DVIPDFM utility

Probably the easiest method for producing acceptable .pdf files is via the utility dvipdfm. This utility is included in distributions of MikTEX, which runs on Windows machines, but it probably needs to be added to your teTEX distribution, if you are running L^{A}TEX on a UNIX machine. The utility and precise information about installing it on your system can be found at the

web page http://gaspra.kettering.edu/dvipdfm/. In essence, this utility

converts a .dvi file into a .pdf file. So, one can first prepare the .dvi file using L^{A}TEX, and then apply the utility dvipdfm to produce the needed .pdf file.^{4} This utility makes inclusion of graphics particularly simple – those that are included in the L^{A}TEX source file are simply converted to the .pdf format. As we note below, things are not so simple with the second alternative, which is to use pdfL^{A}TEX.

pdfL^{A}TEX

An alternative to using dvipdfm to produce .pdf files is to process the source file with pdfL^{A}TEX. This format is available from the standard CTAN sites http://www.ctan.org. It appears that pdfL^{A}TEX and hyperref have some problems when used together. It is necessary to use pdfL^{A}TEX version 14d or later in order to minimize these issues. If your system has an earlier version (most teTEX distributions have version 13d), then you can update your system by retrieving the latest version of pdfL^{A}TEX from

ftp://ftp.cstug.cz/pub/tex/local/cstug/thanh/pdftex/. Even if the

recent versions are used, pdfL^{A}TEX has the same dealing with references embedded with the frontmatter section described above for L^{A}TEX.

4 Beware! The utility dvipdf does not produce acceptable .pdf files, and should not be used. Only dvipdfm should be used to produce .pdf files.

There is one aspect of using pdfL^{A}TEX that bears mention. Many authors want to include EPS^{5} files within their papers. While this is fairly straightforward with L^{A}TEX, there are a couple of points to note when attempting this with pdfL^{A}TEX.

To include a PostScript image in a .pdf file produced with pdfL^{A}TEX, you first have to convert the image to a .pdf file, and then it can be included using the same command sequence as above. The conversion can be accomplished most easily using Ghostscript; you can simply view the file in Ghostview and then print the image to a .pdf file using the pdfwriter option within Ghostview. The result for a standard chess board that is part of the Ghostview distribution is the following image:

Here as well is a copy of a color image. While pdfL^{A}TEX can handle image files in other formats, L^{A}TEX can only handle .eps images reliably.

5 EPS stands for embedded PostScript, which affords a mechanism for including preprepared PostScript files within a L^{A}TEX document.

It also should be noted that you need to have two separate source files – one for L^{A}TEX and one for pdfL^{A}TEX – only in case you wish to insert graphics images in your final paper. If your paper does not include such images, then the same source file can be formatted by both L^{A}TEX and by pdfL^{A}TEX.

Some versions of xdvi can view PostScript files, but often it is necessary to convert the .dvi file to a PostScript file in order to view such images. Even for those versions of xdvi that support embedded PostScript images, the files themselves are binary, and hence can suffer corruption when transferred over the net. Moreover, Ghostview does not support active links at this time, so the advantages of the hyperref package are lost in this case. This explains why we are moving to .pdf files for ENTCS.

5 Summary

The ENTCS macro package should be easy to use and should provide a uniform layout for all the papers in the Proceedings.

Problem 5.1 Finish your paper and get it to your Program Chairman on time!

When you have finished preparing your paper, send a copy of the source file, together with any macro files that are needed to your Program Chairman. If the files are extensive, you can place copies in the pub/incoming sub-directory of the ftp directory on the machine indicated by your Program Chairman using anonymous ftp. If you do this, please send me email to alert me that the file(s) are here.

6 Bibliographical references

ENTCS employs the plain style of bibliographic references in which references are listed in alphabetical order, according the the first author’s last name, and are sequentially numbered. Please utilize this style. We have a BibTEX style file, for those who wish to use it. It is the file entcs.bst which can be found in the directory containing the macro files for ENTCS. The basic rules we have employed are the following:

• Authors’ names should be listed in alphabetical order, with the first author’s last name being the first listing, followed by the author’s initials or first name, and with the other authors names listed as first name, last name.

• Titles of articles in journals should be in emphasized type.

• Titles of books, monographs, etc. should be in quotations.

• Journal names should be in plain roman type.

• Journal volume numbers should be in boldface type, with the year of publication immediately following in roman type, and enclosed in parentheses.

• References to URLs on the net should be “active” and the URL itself should be in typewriter font.

• Articles should include page numbers.

These are illustrated in the following.

References

[1] Civin, P., and B. Yood, Involutions on Banach algebras, Pacific J. Math. 9 (1959), 415–436.

[2] Clifford, A. H., and G. B. Preston, “The Algebraic Theory of Semigroups,” Math. Surveys 7, Amer. Math. Soc., Providence, R.I., 1961.

[3] Freyd, Peter, Peter O’Hearn, John Power, Robert Tennent and Makoto Takeyama, Bireflectivity, Electronic Notes in Theoretical Computer Science 1

   (1995), URL: http://www.elsevier.nl/locate/entcs/volume1.html.

[4] Easdown, D., and W. D. Munn, Trace functions on inverse semigroup algebras, U. of Glasgow, Dept. of Math., preprint 93/52.

[5] Roscoe, A. W., “The Theory and Practice of Concurrency,” Prentice Hall Series in Computer Science, Prentice Hall Publishers, London, New York (1198), 565pp. With associated web site http://www.comlab.ox.ac.uk/oucl/

   publications/books/concurrency/.

[6] Shehadah, A. A., “Embedding theorems for semigroups with involution, “ Ph.D. thesis, Purdue University, Indiana, 1982.

[7] Weyl, H., “The Classical Groups,” 2nd Ed., Princeton U. Press, Princeton, N.J., 1946.